Camino Company manufactures designer to-go coffee cups. Each line of coffee cups is endorsed by a high- profile celebrity and designed with special elements selected by the celebrity. During the most recent year Camino Company had the following operating results while operating at 75 percent (75.000 units) of its capacity Sales revenue Cost of goods sold $ 900,000 375,000 Gross profit Operating expenses S525,000 18,750 Net operating incomeS 506.250 Caminos cost of goods sold and operating expenses are 80 percent variable and 20 percent fixed. Camino has received an offer from a professional wrestling association to design a coffee cup endorsed by its biggest star and produce 20,000 cups for $8 each (total $160,000). These cups would be sold at wrestling matches throughout the United States. Acceptance of the order would require a $60,000 endorsement fee to the wrestling star, but no other increases in fixed operating expenses. Required: 1. Complete the incremental analysis of the special order in the table provided below Per cup Total Incremental Revenue Less: Incremental Costs Variable COGS Variable Operating Expenses Endorsement Fee Incremental Profit
media%2F847%2F847012ca-3a8e-4468-968a-83
Answer #1

(1) Incremental Analysis of the special order :-

                

Per Cup

Total (20000 Units)

Incremental Revenue

8

(20000*8) = 160000

Less:- Incremental cost

Variable COGS

(375000/75000)*80% = 4

(20000*4) = 80000

Variable Operating Exp

(18750/75000)*80% =0.2

(0.2*20000) = 4000

Endorsement Fee

(60000/20000) = 3

60000

Incremental Profit

0.8

16000

(2) YES, Camino should accept the order

(3) If camino operating at full capacity, Price per cup:-

                  900000/75000 = 12